# Continual Forgetting for Pre-trained Vision Models

### v1

@@ 2.1 Continual Learning
Continual learning aims to enable models to acquire new knowledge without forgetting previously learned information [35]. It is a learning paradigm that is particularly applicable in dynamic and changing scenarios. Researchers have designed three strategies to achieve this goal, including rehearsal-based methods [37, 57, 30, 46, 59, 64, 85, 86], regularization-based methods [35, 41, 2, 61, 84], and structure-based methods [49, 79, 60, 16, 1, 45]. These three strategies for continual learning are frequently combined to improve performance [57, 46, 87].
@@ 2.3 Parameter-Efficient Fine-Tuning
Training large models by self-supervised learning and then fine-tuning them on downstream tasks has become a new paradigm of deep learning [7, 54, 27, 26, 55, 39, 71, 70, 69, 51]. Parameter-efficient fine-tuning (PEFT) techniques [29, 40, 28, 33, 11, 83, 81] are proposed to optimize a limited number of parameters, as fully fine-tuning increasing large models [54, 7, 34] becomes less practical for various downstream tasks.
@@ 4.1 Overview
In this study, we propose the Group Sparse LoRA (GS-LoRA) to achieve parameter-efficient and effective continual forgetting. Fig. 2 shows the overall pipeline of GS-LoRA. We consider each forgetting request as a task in the continual forgetting process. For each task $\mathcal{T}_{t}$ where $t=1,2,\cdots,T$, we use LoRA to fine-tune the FFN modules in Transformer blocks. To mitigate catastrophic forgetting of the remaining knowledge, smaller network changes are preferred [48, 72, 79, 49]. Therefore, we use group sparse regularization to select and modify fewer blocks. To achieve the optimization goal in Eq. 3, we use a selective forgetting loss to maximize the original loss in the forgotten classes and a knowledge retention loss to minimize the loss in the remaining classes.
@@ 4.2 GS-LoRA
Group Sparsity Selection. To mitigate catastrophic forgetting and achieve precise modifications automatically, we introduce a group sparsity selection strategy that enables the selection of fewer Transformer blocks. We utilize group Lasso which can efficiently select parameters for certain groups [78, 73, 18, 43] and zero out others. Suppose LoRA matrices added to the $\ell$-th Transformer block in task $\mathcal{T}_{t}$ are ${\mathbf{B}_{1}}^{(\ell)}_{t},{\mathbf{A}_{1}}^{(\ell)}_{t},{\mathbf{B}_{2}}^{(\ell)}_{t},{\mathbf{A}_{2}}^{(\ell)}_{t}$. Then the optimization goal with group sparse regularization can be expressed as follows:
@@ 4.3 Loss Function
Table 1: Single-step forgetting results for face recognition. $Acc_{r}$ and $Acc_{f}$ are the accuracies of remaining and forgotten classes. $*{}^{*}$ denotes the original methods with a rehearsal buffer. Note that “retrain” represents retraining the model using replay data and the training epoch is the same as other methods to ensure a fair comparison. Pre-train denotes the results before forgetting. All setting is in the form of 100-Y, which means all experiments start from a pre-trained model (100 classes originally) and forget Y classes.
@@ 5.1 Experimental Setup
Datasets and Pre-trained Models. We evaluate the effectiveness and efficiency of GS-LoRA using published Transformer-based models in a privacy-sensitive face recognition task and a general object detection task. More experiments on image classification can be found in the Supplementary Material. For the face recognition task, we constructed a subdataset called CASIA-Face100 which collects 100 face IDs from the CASIA-WebFace [77] dataset. We use a Face Transformer [82] with six Transformer blocks pre-trained on the CASIA-Face100 dataset. For the object detection task, we use deformable DETR [88] which has six encoders and six decoders pre-trained on the COCO 2017 [42] dataset.
@@
Here $Acc_{r}^{(t)}$ is calculated on the retained dataset after task $\mathcal{T}_{t}$ and $Drop^{(t)}=Acc_{f}^{(t-1)}-Acc_{f}^{(t)}$ is the performance drop on forgotten classes before and after training the taskWe take the classification problem as an example to define $H\text{-}Mean$. Replace $Acc$ with $AP$ for object detection tasks.. After learning task $\mathcal{T}_{t}$, we evaluate the performance on all previously forgotten classes in task $\mathcal{T}_{1},\mathcal{T}_{2},\cdots,\mathcal{T}_{t-1}$.
@@ 6.2 Scalability
6.3 Different Grouping Strategies
We consider three grouping strategies: “Block”, “Module” and “Matrix”. “Block” is the default setting. “Module” denotes each LoRA module is a group, resulting in twice the number of groups compared to the Transformer blocks. “Matrix” means each matrix in LoRA modules is a group and the number of groups is four times the number of Transformer blocks.
We conduct our experiments on a Face Transformer and all the experiments are performed at the same sparse intensity, i.e. $\alpha$ is the same. Tab. 8 shows the results of different grouping strategies. Notably, all grouping strategies yield exceptional performance. It is observed that with a more detailed grouping strategy, the zero-group ratio increases, which makes sense because each group has fewer parameters and more flexibility.
@@ 6.3 Different Grouping Strategies
Grouping
Strategy
$Acc_{f}\downarrow$ $Acc_{r}\uparrow$ $H\uparrow$
Zero Group
Ratio
Block 1.97 71.06 71.43 0.17
Module 0.93 70.58 71.70 0.50
Matrix 1.51 70.36 71.30 0.58
Table 8: Effect of grouping strategies. The zero-group ratio goes up when a more detailed grouping strategy is used.
@@ B.1 Face Recognition
Forgetting. For the forgetting process, we use 1337 as the random seed to generate a forgetting order. Implementation details can be found in Tab. S2 for all experimental settings, which shows the robustness of GS-LoRA. Here, BND is the bound in Eq. (10), K and $\alpha_{K}$ is the hyperparameter in Eq. (9), and $\beta$ is the hyperparameter in Eq. (12). Note that “data ratio” is the ratio of data used for forgetting to data used for pre-training. To achieve fast model editing, the forgetting epoch is set to 100 and 0.1$\times$ pre-training data is used. This is equivalent to fine-tuning the model using all pre-training data with only 10 epochs, which is less than 1% compared with 1200 epochs in the pre-training process.
Figure S3: Visualization for continual forgetting. The left column shows the results from the pre-trained model. The middle column shows the results when “keyboard” (red bounding boxes in the left column) is erased. The left column shows the results when more objects (e.g., person, book, chair) are erased.
@@ B.3.1 Continual Learning Methods
We implement six continual learning methods to realize continual forgetting. Taking EWC as an example, we conduct it for forgetting as follows. First, we give randomly wrong labels to the forget set. Then, we use the remaining set to calculate the weight importance of EWC. Finally, we regard learning on the modified forget dataset as a new task and perform EWC algorithm. Note that we freeze the final FFN layer to ensure backbone forgetting. Additionally, for a fair comparison, we use the remaining set as a replay buffer to enhance the performance, which is denoted as $*{}^{*}$.
@@ C.2 Ablations on $\beta$ in Loss Function
Table S6: Experiment results with incomplete replay. Thirty classes are forgotten in all experiments. In the remaining 70 classes, only some classes can be replayed. Each subtable shows the results with a different number of replay classes. Pre-train denotes the results before forgetting. $*{}^{*}$, $*{}^{*}$ and $*{}^{*}$ denote the original methods with a rehearsal buffer. LoRA denotes using LoRA to fine-tune FFN modules in Transformer blocks without group sparse. Our method (GS-LoRA) is highlighted in color. We specifically evaluate the accuracy of the classes without replay samples and report it as $Acc_{r}^{\dagger}$.
C.3 More Experiments with Incomplete Replay

### v2

@@ 2.1 Continual Learning
Continual learning aims to enable models to acquire new knowledge without forgetting previously learned information [37]. It is a learning paradigm that is particularly applicable in dynamic and changing scenarios. Researchers have designed three strategies to achieve this goal, including rehearsal-based methods [39, 60, 32, 48, 63, 69, 93, 94, 11], regularization-based methods [37, 43, 2, 65, 92], and structure-based methods [51, 87, 64, 18, 1, 47]. These three strategies for continual learning are frequently combined to improve performance [60, 48, 95, 54].
@@ 2.3 Parameter-Efficient Fine-Tuning
Training large models by self-supervised learning and then fine-tuning them on downstream tasks has become a new paradigm of deep learning [7, 57, 29, 28, 58, 41, 77, 76, 75, 53]. Parameter-efficient fine-tuning (PEFT) techniques [31, 42, 30, 35, 12, 91, 89] are proposed to optimize a limited number of parameters, as fully fine-tuning increasing large models [57, 7, 36, 78] becomes less practical for various downstream tasks.
@@ 4.1 Overview
Considering two key challenges in Sec. 1 and the optimization goal in Eq. 3, we propose Group Sparse LoRA (GS-LoRA) with selective forgetting loss and knowledge retention loss to achieve continual forgetting. Fig. 2 shows the overall pipeline of GS-LoRA. To achieve efficient forgetting, we use LoRA to fine-tune the FFN modules in Transformer blocks. To mitigate catastrophic forgetting of the remaining knowledge, smaller network changes are preferred [50, 79, 87, 51]. Therefore, we use group sparse regularization to select and modify fewer blocks. Sec. 4.2 gives a more detailed description. To achieve the optimization goal in Eq. 3, we use a selective forgetting loss to maximize the original loss for the forgotten classes and a knowledge retention loss to minimize the loss for the remaining classes in Sec. 4.3.
@@ 4.2 GS-LoRA
Group Sparsity Selection. To mitigate catastrophic forgetting and achieve precise modifications automatically, we introduce a group sparsity selection strategy that enables the selection of fewer Transformer blocks. Although there are many ways to conduct a selection like routers [86, 67], meta learning [73], neural architecture search [62, 14], we utilize group Lasso, known for its simplicity and effectiveness in selecting parameters for specific groups [85, 80, 20, 45] while setting others to zero. Suppose LoRA matrices added to the $\ell$-th Transformer block in task $\mathcal{T}_{t}$ are ${\mathbf{B}_{1}}^{(\ell)}_{t},{\mathbf{A}_{1}}^{(\ell)}_{t},{\mathbf{B}_{2}}^{(\ell)}_{t},{\mathbf{A}_{2}}^{(\ell)}_{t}$. Then the optimization goal with group sparse regularization can be expressed as follows:
@@ 4.3 Loss Function
Table 1: Single-step forgetting results for face recognition. $Acc_{r}$ and $Acc_{f}$ are the accuracies of remaining and forgotten classes.  denotes the original methods with a rehearsal buffer. Note that “retrain” represents retraining the model using replay data and the training epoch is the same as other methods to ensure a fair comparison. Pre-train denotes the results before forgetting. All setting is in the form of 100-Y, which means all experiments start from a pre-trained model (100 classes originally) and forget Y classes.
@@ 5.1 Experimental Setup
Datasets and Pre-trained Models. We evaluate the effectiveness and efficiency of GS-LoRA using published Transformer-based models in face recognition tasks and object detection tasks. More experiments on image classification can be found in the Supplementary Material. For the face recognition task, we constructed a subdataset called CASIA-Face100 which collects 100 face IDs from the CASIA-WebFace [84] dataset. We use a Face Transformer [90] pre-trained on the CASIA-Face100 dataset. For the object detection task, we use a deformable DETR [96] pre-trained on the COCO 2017 [44] dataset.
@@
Here $Acc_{r}^{(t)}$ is calculated on the retained dataset after task $\mathcal{T}_{t}$ and $Drop^{(t)}=Acc_{f}^{(t-1)}-Acc_{f}^{(t)}$ is the performance drop on forgotten classes before and after training the task. [^1] We take the classification problem as an example to define $H\text{-}Mean$. Replace $Acc$ with $AP$ for object detection tasks. After learning task $\mathcal{T}_{t}$, we evaluate the performance on all previously forgotten classes in task $\mathcal{T}_{1},\mathcal{T}_{2},\cdots,\mathcal{T}_{t-1}$.
@@ B.1 Face Recognition
Forgetting. For the forgetting process, we use 1337 as the random seed to generate a forgetting order. Implementation details can be found in Tab. S2 for all experimental settings, which shows the robustness of GS-LoRA. Here, BND is the bound in Eq. (10), K and $\alpha_{K}$ is the hyperparameter in Eq. (9), and $\beta$ is the hyperparameter in Eq. (12). Note that “data ratio" is the ratio of data used for forgetting to data used for pre-training. To achieve fast model editing, the forgetting epoch is set to 100 and 0.1$\times$ pre-training data is used. This is equivalent to fine-tuning the model using all pre-training data with only 10 epochs, which is less than 1% compared with 1200 epochs in the pre-training process.
Figure S3: Visualization for continual forgetting. The left column shows the results from the pre-trained model. The middle column shows the results when “keyboard" (red bounding boxes in the left column) is erased. The left column shows the results when more objects (e.g., person, book, chair) are erased.
@@ B.3.1 Continual Learning Methods
We implement six continual learning methods to realize continual forgetting. Taking EWC as an example, we conduct it for forgetting as follows. First, we give randomly wrong labels to the forget set. Then, we use the remaining set to calculate the weight importance of EWC. Finally, we regard learning on the modified forget dataset as a new task and perform EWC algorithm. Note that we freeze the final FFN layer to ensure backbone forgetting. Additionally, for a fair comparison, we use the remaining set as a replay buffer to enhance the performance, which is denoted as EWC.
@@ C.2 Ablations on $\beta$ in Loss Function
Grouping
Strategy
$Acc_{f}\downarrow$ $Acc_{r}\uparrow$ $H\uparrow$
Zero Group
Ratio
Block 1.97 71.06 71.43 0.17
Module 0.93 70.58 71.70 0.50
Matrix 1.51 70.36 71.30 0.58
Table S6: Effect of grouping strategies. The zero-group ratio goes up when a more detailed grouping strategy is used.
C.3 Different Grouping Strategies
We consider three grouping strategies: “Block", “Module" and “Matrix". “Block" is the default setting. “Module" denotes each LoRA module is a group, resulting in twice the number of groups compared to the Transformer blocks. “Matrix" means each matrix in LoRA modules is a group and the number of groups is four times the number of Transformer blocks.
@@ C.3 Different Grouping Strategies
Table S7: Experiment results with incomplete replay. Thirty classes are forgotten in all experiments. In the remaining 70 classes, only some classes can be replayed. Each subtable shows the results with a different number of replay classes. Pre-train denotes the results before forgetting. L2, MAS and EWC denote the original methods with a rehearsal buffer. LoRA denotes using LoRA to fine-tune FFN modules in Transformer blocks without group sparse. Our method (GS-LoRA) is highlighted in color. We specifically evaluate the accuracy of the classes without replay samples and report it as $Acc_{r}^{\dagger}$.
We conduct our experiments on a Face Transformer and all the experiments are performed at the same sparse intensity, i.e. $\alpha$ is the same. Tab. S6 shows the results of different grouping strategies. Notably, all grouping strategies yield exceptional performance. It is observed that with a more detailed grouping strategy, the zero-group ratio increases, which makes sense because each group has fewer parameters and more flexibility.
C.4 More Experiments with Incomplete Replay